### v1

# Defining a Role-Centered Terminology for Physical Representations and Controls

Abstract.
Previous classifications advanced research through a better understanding of the field and the variety of tangible user interfaces and related physical user interfaces, especially by discretizing a degree of tangibility based on the specimens produced by the community over the years, since the conceptualization of Tangible User Interface initiated a research effort to deepen the exploration of the concept. However, no taxonomy enables the classification of tangible user interfaces at the application level. This article proposes to refine the description of tangible user interfaces’ interactional components through a terminological approach. The resulting terms are blended words, built from known words, that self-contain what digital role is represented or controlled and how it becomes physical. This holistic terminology then enables the definition of applications’ hallmarks and four classes of tangibility for applications, which surpass the description of physical user interface specimens’ morphology by abstracting and discriminating specimens at the applicative level. The descriptiveness and holisticness of the new terminology, as well as the clustering and discriminative power of the limited number of four classes, are showed on a corpus of applicative tangible user interfaces’ specimens from the literature. Promising future work will benefit from the holistic terminology, the applications’ hallmarks, and the tangibility classes, to describe applicative tangible user interfaces and related physical user interfaces to better understand the dozens of specimens that were produced by the field over three decades. Indeed, describing and classifying this whole set would deepen our understanding to provide tools for future developers and designers.
terminology, linguistics, blended words, physical computing, physical user interfaces, graspable user interfaces, tangible user interfaces, embodied interfaces, embedded interfaces, taxonomy, classification, role-centered viewpoint, what–how terminology, tangibility classes of applications
$c$s: Human-centered computing HCI theory, concepts and modelscopyright: ccjournalyear: 2025conference: arXiv.org, Cornell University; 2025; Ithaca, NY, USA
1. Introduction
Conceptualization often precedes a community’s focus on the same research path. Whereas precursor specimens—of what was later conceptualized as Tangible User Interfaces (TUIs)—arose in the mid-1970s and early 1980s Shaer and Hornecker, 2010: Ch. 2, Ullmer et al., 2022, physical computing came to the forefront of Human–Computer Interaction (HCI) researchers’ attention only fifteen years later (Wellner, 1991; Hinckley et al., 1994; Fitzmaurice et al., 1995), after the concept of ubiquitous computing was conceptualized in the late 1980s (Weiser, 1991, 1999). Several works initiated physical computing’s exploration, from “passive real-world props” (Hinckley et al., 1994), to “graspable user interfaces” (Fitzmaurice et al., 1995), to “manipulative user interfaces” (Harrison et al., 1998), to “embodied user interfaces” (Fishkin et al., 1999), which were the growing roots of the same tree (Fishkin, 2004). Likewise, the deeper exploration of the design and technology of tangible user interfaces began after they were conceptualized in 1997 (Ullmer and Ishii, 1997). Since then, while the number of specimens has grown yearly, taxonomies, paradigms, and frameworks have emerged Hoven and Eggen, 2004, Mazalek and Hoven, 2009, Shaer and Hornecker, 2010: Ch. 5, defining terms to describe the roles of physical artifacts (e.g., phicon (Ullmer and Ishii, 1997), token (Holmquist et al., 1999; Shaer et al., 2004; Ullmer et al., 2005), and tangible representation (Ullmer and Ishii, 2000)). However, no global terminology is shared in the literature, even if refining and conceptualizing tangible user interfaces’ interactive components better could help refine the design space, deepen specimens’ exploration, and broaden applications’ possibilities—indeed, sharing a common vocabulary and the same concepts also helps aggregating research efforts on a common path.
Previous work provided isolated terms, common terms, and general terms, but none can distinguish all the entities involved in tangible user interfaces in a single namespace. On the one hand, isolated terms are limited in the scopes and paradigms that define them (e.g., Token And Constraints (Shaer et al., 2004)), and, therefore, remain separated in diverse namespaces. On the other hand, some terms are found across several research works, but with inconsistent meanings (e.g., “tokens” (Holmquist et al., 1999; Shaer et al., 2004; Ullmer et al., 2005)). As a consequence, many researches are built upon general terms (e.g., “tangibles,” “physical object,” and “tangible object” (Shaer and Hornecker, 2010)), thus escaping terminological ambiguity and describing all physical artifacts as a whole, without distinction. However, such general terms remain broad and undescriptive. For instance, the term “tangibles” is plural and cannot name merely a single artifact. The terms “physical object” and “tangible object” lend a new meaning to the term “object”, which is already overused in computer science (Shaer et al., 2004). Finally, “objects” are already “physical” and “tangible” in the usual sense.
Furthermore, scopes other than pure tangible user interfaces resort to physical entities as representations of controls, but not especially for data representation (e.g., in Augmented Reality (Cordeil et al., 2020) and Virtual Reality (Mahieux et al., 2022) scopes). The lack of precise terms can lead to consider any user interface resorting to physical objects as a “tangible user interface,” even if these objects are only a subpart of a user interface that does not represent data with physical bodiement, thus leading to confusion and misunderstanding (e.g., before general frameworks, recognizing tangible user interfaces was mainly relying on “I know one when I see one” approaches (Fishkin, 2004)). Fishkin’s taxonomy overcame these considerations by “treating tangibility as a spectrum rather than a binary quantity” (i.e., “tangible” and “not tangible” user interfaces) and organizing “proof of concept” specimens over a twenty-item discretized space (Fishkin, 2004). However, this taxonomy, which has contributed to better delineating and understanding the field, does not address tangible user interfaces at the applicative level (i.e., user interfaces that gather and assemble components to meet the needs of final users). Indeed, this taxonomy is mainly focused on analyzing data representations (i.e., distance between data output and the inputs enabling data manipulation and edition, and the user actions’ degree of analogy to perceive and manipulate data) and describes user interfaces as wholes (i.e., without refining the components that compose applicative user interfaces). Except that such a global description requires determining what is “the” datum of the user interface, but this often becomes tricky in applicative user interfaces, which involve multiple tasks, outputs, or users. Although describing the morphology and meaning of tangible user interfaces globally enabled refining the concept of “tangibility” in physical user interfaces, developing applications of such user interfaces may require describing components separately and beyond data representations.
For instance, the Urp specimen (Urban Planning Workbench) melts, on the surface of a table, models of buildings with editing and navigation tools and with measurement objects, and all of these are augmented with video-projection to display information from the computational simulation.
For instance, the Urban Planning Workbench (Urp) (Ben-Joseph et al., 2001; Ishii et al., 2002; Underkoffler and Ishii, 1999) specimen enables the manipulation of building models on a surface table, along with editing and navigation tools, as well as measurement objects. All of these features are augmented with video projection to display information from the computational simulation.
Describing this “workbench” application globally, as a whole, ignores all the subtleties of its functioning and, for example, the role played by intangible representations. As well, determining the distance between input and output fails when an applicative tangible user interface only conveys data.
In addition, terms describing intangible representations are underrepresented: all the previous terms—isolated, common, and general ones—focused nearly exclusively on detailing tangible representations (whereas many terms were defined for GUIs, such as icons, windows, or sliders). Consequently, could we discriminate, in a single namespace, all the tangible, graspable, and intangible entities found in tangible user interfaces (which represent data physically) and in any other related physical user interfaces? Since previous work provided a profusion of specimens, could we use the refined terminology to better discriminate tangible user interfaces according to the description of their components? Consequently, could such a refined terminology of physical entities lead to a classification of applicative tangible user interfaces?
This article proposes a role-centered terminology that names physical entities by defining twelve terms that self-contain what is represented or controlled and how. This terminology serves two objectives: the first is to provide terms that better align with foundational definitions, thereby improving conceptual understanding; the second is to refine the description of the representations and controls involved in tangible and related physical user interfaces. To show the descriptive power of this terminology, this article successfully discriminates all the physical representations and controls found in a collection of applicative tangible user interfaces. Furthermore, this holistic discrimination also enables the classification of applicative tangible user interfaces into four tangibility classes. We envision and expect that such terminological and classification advances will serve as a foundation for future work to investigate the composition of previous applicative tangible user interfaces and related physical user interfaces in greater depth and to provide tools and toolkits for future designers.
After reviewing related work, this article introduces a new terminology based on a taxonomy with two axes, and then derives four tangibility classes of applicative tangible user interfaces from this terminology. Subsequently, the descriptiveness of the terminology is validated on a corpus of applicative tangible user interfaces, which are then clustered into the four classes. Then, previous terms from the literature are mapped to the new terminology, and the four classes are linked to previous classifications. Finally, before concluding, some limitations are highlighted, and directions are provided for future work.
2. Related Work
This section reviews related work on terms for physical representations and controls in tangible user interfaces, the creation of new words in the usual language through the word blending, and taxonomies and classifications of tangible user interfaces.
2.1. Terms for Physical Representations
The term “tangible” appeared early to describe a “tangible model” for 3D input in Computer-Aided Architectural Design (CAAD) (Aish and Noakes, 1984) and “tangible manipulation” in the Digital Desk (Wellner, 1991) (but the latter was later renamed “tactile interaction” (Wellner, 1993)). Later, the terms “graspable” and “tangible” were introduced to define Graspable User Interfaces (Fitzmaurice et al., 1995) and Tangible User Interfaces (Ullmer and Ishii, 1997). Finally, the term “intangible” appeared to name audio and video output modalities in the MCRit model [^1] MCRit: “Model-Control-Representation (intangible and tangible)” (Ullmer, 2002; Ullmer et al., 2005), first known as MCRpd: “Model-Control-Representation (physical and digital)” (Ullmer and Ishii, 2000). (Ullmer, 2002; Ullmer et al., 2005). Whereas the term “tangible” is used close to dictionary definitions, the term “intangible” differs from the usual sense: it simply means “not tangible.” Therefore, intangible representations are those generated by the computer; they may disappear when the computer is turned off (e.g., sound, video projection, and any audio and visual modalities or displays). This section summarizes the usual terms used in the literature to name the interactional components of tangible user interfaces, categorized into general, isolated, and shared terms, as well as terms employed in specific toolkits and design tools.
2.1.1. General Terms
Some terms are used in the literature to generically name the physical part of tangible user interfaces that embody digital information, including “physical artifacts” (Ullmer and Ishii, 2000), “physical representations” (Ullmer and Ishii, 2000), “tangible representations” (Ullmer, 2002; Ullmer et al., 2005), “tangibles” (Ullmer, 2002), and “tangible objects” (Shaer and Hornecker, 2010). Those terms are variations to name graspable and tangible interactional components broadly. However, the common use of the term “tangible” was initially defined for bodied data representations (i.e., not graspable anchors on tabletops). Finally, only one general term refers to neither graspable nor tangible components: “intangible representations” (Ullmer, 2002; Ullmer et al., 2005).
2.1.2. Isolated Terms
The term passive real-world prop (Hinckley et al., 1994) is used to name bodied representations of tools and data that map sensed user actions directly to control digital data displayed on a distant screen. Props can be symbolic (Ullmer and Ishii, 2000) (e.g., a sphere (Hinckley et al., 1994)) or iconic (Ullmer and Ishii, 2000) (e.g., a doll’s head (Hinckley et al., 1994)) data representations, combined with tools (e.g., a slicing plate (Hinckley et al., 1994)). The term “passive” refers to the system’s ability to track and reproduce movements made on the props. We can understand the choice of this term with the first loop of tangible interaction (Ishii, 2008): props provide a first passive feedback loop before computation’s feedback arises—the relative position of the props informs about cutting plane location without watching the screen (what is reinforced when using the doll’s head (Hinckley, 1996)).
The term physical handle (Fitzmaurice et al., 1995) (or phandle (Ullmer and Ishii, 1997)) refers to physical artifacts of general form (e.g., bricks and pucks) that can be attached to intangible shapes displayed on a surface. Once the physical handle acquires an intangible shape, motions are transmitted to control the digital object (e.g., position, orientation, and size).
The term phicon (Ullmer and Ishii, 1997) stands for “physical icons” of a specialized form that instantiates graphical user interface (GUI) components physically. Once put on a horizontal surface, phicons are linked to digital data (e.g., a building (Ullmer and Ishii, 1997)), and intangible data is updated to match phicons’ physical states (e.g., a displayed map is updated (Ullmer and Ishii, 1997)).
2.1.3. Shared Terms
Some terms do not name the artifacts by what they are but by their role and appear across several research works. For example, the terms “token,” “container,” and “tool” (Holmquist et al., 1999) define three roles in giving access to digital information. Tokens are specialized artifacts whose physical properties match the represented digital information (e.g., phicons (Ullmer and Ishii, 1997)). Containers are generic artifacts that can represent any type of information. Whereas tokens are continuously assigned to the same digital information, containers are associated with any digital information over time. Finally, tools are physical artifacts that have computational functions, enabling them to act on the displayed data (e.g., physical handles (Fitzmaurice et al., 1995)). Beyond naming physical artifacts, access points to digital information are also named: digital information is rendered by information faucets (e.g., display devices, speakers, and tactile devices). However, the “Token+Constraint” approach (Ullmer et al., 2005) suggests that all those artifacts are “tokens”: “containers” are “symbolic tokens with permanent binding,” “tokens” are “iconic tokens with dynamic binding,” and “tools” are “tokens bound to operations” (Ullmer et al., 2005).
The “Token+Constraint” approach (Ullmer et al., 2005) reuses the term “token” in its usual sense, which is also its standard usage in computer science. The term “token” is then used in combination with the term “constraint,” where tokens are physical objects representing digital information, and constraints are physical regions bound to digital operations. Physical constraints structure the placement and combination of tokens, which systems can sense and interpret to trigger operations.
The term token is also used in the usual sense in the “TAC Paradigm” (Shaer et al., 2004), based on the aforementioned “Token+Constraint” approach (Ullmer, 2002; Ullmer et al., 2005). However, the TAC Paradigm first names physical objects composing a tangible user interface under the term pyfo to move away from the various uses of the term “object.” Pyfo is a constructive and recursive term (not atomic): pyfos are made of pyfos (e.g., the six sides of a box are also pyfos (Shaer et al., 2004)). As well, pyfos are not limited to whole graspable physical objects: pyfos are also intangible physical objects (e.g., intangible constraints drawn on a surface, such as Senseboard’s displayed grid, are pyfos (Jacob et al., 2002)). Thus, tokens are graspable pyfos representing digital information or computational functions; constraints are graspable or intangible pyfos that structure the behavior of tokens. Beyond physical representations, TAC Paradigm’s terminology also includes terms for digital elements, such as the term variable, which names digital information or computational functions. Whereas some variables can be bound to tokens, semantic variables cannot. Finally, systems behave according to a series of rules: association triplets between a token, a variable, and some constraints. These rules are referred to as TAC (i.e., Token And Constraints).
2.1.4. Terms in Toolkits and Design Tools
Toolkits and design tools also require terms to assign names to physical representations and controls, as well as to differentiate their roles.
The term phidgets (Greenberg and Fitchett, 2001) stands for “physical widgets”: packaged, low-level devices and software architecture inspired by GUI toolkits to ease the building of physical user interfaces. The Phidgets toolkit provides physical primitives for controllers, sensors, and mechanical parts that are commercially available [^2] Phidgets’ website: https://www.phidgets.com/, last accessed 2022-06-09..
ASUR++ (Dubois et al., 2002, 2003) is a notation for describing the physical and digital entities of mobile mixed systems. This notation extends the ASUR notation for mixed reality systems and is usable for tangible user interfaces (Dubois et al., 2003). The notation defines two terms for physical objects, by distinction of two digital roles: tools that support a task (Rtool) and targeted objects that are modified by a task (Robject).
The ROSS API (Wu et al., 2012) defines four nested levels to describe physical representations and controls. A strength of this low-level framework is that it goes beyond the description of physical objects (RObject) by denoting active spaces (RSpace) and surfaces (RSurface) that can sense interaction. However, even if input controls—such as buttons, switches, and sliders—are described (RControl), physical objects’ roles—such as tools and data representations—are not describable by the API.
2.1.5. Existent Terms’ Limitations
The previous terms are built from two main strategies: drawing a parallel with GUI elements (e.g., “icon,” “widget,” “handle,” and “container”) or naming their physical roles (e.g., “prop,” “token,” and “constraint”). At certain points, these approaches enable the description of low embodiment (e.g., “phandles”) and higher embodiment (e.g., “phicon”). However, these approaches fail to establish a global terminology that can describe all the components residing within tangible user interfaces. For example, “tokens” can equally name artifacts representing data or functions. Moreover, tangible user interfaces often comprise representations that are not “purely tangible” (i.e., not representing data), but no terms distinguish a pure tangible user interface from other related physical user interfaces. Thereby, the lack of descriptive and discriminative power leads to the wide use of general terms to name any artifact indistinctly.
Some of the terms listed above come from everyday objects (e.g., “tokens,” “props,” “handles,” “faucets,” “bricks,” or “pucks”); some are blended words created specifically (e.g., “phandle,” “phicon,” or “phidget”); and some others associate two words (e.g., “physical artifact,” “tangible representation,” or “tangible object”). Most terms describe the artifacts but not how they relate to the digital world, which is suggested by a metaphor of their original use transposed to the digital world (e.g., “tokens,” “props,” and “faucets”). Only a few support the idea of digital entities coming to the physical world (such as “phidgets” and “phicons”)—but they are isolated terms—and only a few express the coupling with a digital entity (e.g., “tangible representation” and “physical representation”).
As a general term, this article considers that “physical representations” comprise tangible, graspable, and intangible representations (as pyfos that also consider intangible representations as being physical)—the difference between “tangible” and “graspable” representations residing in the degree of embodiment: high and low, respectively. High embodiment occurs when some of the digital entities’ representational characteristics are embedded in their representations (similarly to the specialized form artifact definition (Fitzmaurice and Buxton, 1997)) and have iconic meaning (Ullmer and Ishii, 2000). Low embodiment occurs when no digital entities’ representational characteristics are embodied by their representations (similarly to the generic form artifact’s definition (Fitzmaurice and Buxton, 1997)) and that have symbolic meaning (Ullmer and Ishii, 2000). This article also considers that the term “artifacts” refers to bodied objects, regardless of whether they are graspable or tangible. This article introduces a new holistic terminology that refines the descriptive level of the aforementioned physical representations by employing word-blending formation processes.
2.2. Blended Words
Linguistics studies the evolution of everyday language over time, including word-formation processes (e.g., derivations, abbreviations, and acronyms). Some new words are created by combining splinters (Barrena Jurado, 2019) excerpted from source words. For example, blended words concatenate a basis from a first word and a suffix from a second word, with phonemic overlap (Beliaeva, 2016; Connolly, 2013; Gries, 2004). Linguists observed several possibilities and strategies of splinters excerption and combination that enter the definition of “blended words” that keep source words recognisable (Gries, 2004, 2006). For instance, “heliport” (heli(copter) + (air)port), “motel” (mot(or) + (h)otel), and “brunch” (br(eakfast) + (l)unch) are blend words (Connolly, 2013; Gries, 2004). However, blends can follow some other word-formation processes. For example, clipping compounds concatenate the bases of two words (Beliaeva, 2016; Gries, 2004, 2006) (e.g., “fintech” (fin(ancial) tech(nology)), “sysadmin” (sys(tem) + admin(istrator)), and “hydrail” (hyd(rogen) + rail(way))). Word blending enables creating new words of only 1, 2, or 3+ syllables (Connolly, 2013), thus creating new vocabulary that designates concepts through single words, which also facilitates communication.
Blended words already exist in Computer Science and Human–Computer Interaction: for example, “pixel” (pic(ture)s + el(ement)) and “widget” (wi(ndow) + (ga)dget) are kinds of blended words that have become common in the usual language and are referenced in standard dictionaries. In the field of tangible user interfaces, several of the aforementioned terms are blended words, such as “phicon” (ph(ysical) + icon) (Ullmer and Ishii, 1997), “phandle” (ph(ysical) + (h)andle) (Ullmer and Ishii, 1997), and “phidget” (ph(ysical) + (w)idget) (Greenberg and Fitchett, 2001).
This article employs word blend strategies to create new three-syllable terms that name physical entities found in tangible and related physical user interfaces.
2.3. Taxonomies and Classifications
The literature has resorted to taxonomies and classifications to characterize the conceptual space of physical user interfaces. Whereas the seminal work on graspable user interfaces proposed a conceptual space with thirteen axes (Fitzmaurice et al., 1995), Ishii later classified tangible user interfaces through eight genres (Ishii, 2008). About two dozen frameworks organized the conceptual space along different numbers of axes Mazalek and Hoven, 2009, Shaer and Hornecker, 2010: Ch. 5, Ullmer et al., 2022: Ch. 3.
For example, Fishkin’s taxonomy of physicality (or tangibility) (Fishkin, 2004) organized physical user interfaces as a twenty-item conceptual space along two axes: embodiment (i.e., the physical relationship between the input and the output of interaction loops) and metaphor (i.e., the digital meaning given to interaction loops with physical representations).
All those frameworks were organized by types: abstracting, designing, and building; and by facets: experiences, domains, physicality, interactions, and technologies (Mazalek and Hoven, 2009).
This article proposes a new taxonomy on two axes for the physical entities found in applicative tangible user interfaces and related physical user interfaces, based on the role (i.e., what) and tangibility (i.e., how) of the digital entities that are represented or controlled. A terminology is then built by taking the values of these two axes as source words for a blended-word formation process.
3. Proposing a Terminology and a Classification
This section introduces a new terminology that shifts the existing viewpoint from “physical artifacts that are coupled to the digital world” to a role-centered viewpoint of “data coming to the physical world.” Before defining the terminology, the definition of physical entities is refined, and then the concept of a tangible user interface is recalled. Finally, an abstraction of applicative tangible user interfaces and related physical user interfaces is proposed, along with their classification into four tangibility classes.
3.1. Physical Interactional Entities
Let us first define the set of physical interactional entities involved in tangible and related physical user interfaces, which the new terminology aims to name. Indeed, the definitions in this article build upon the definition of “Physical Interactional Entities,” which are considered the highest level of abstraction of anything that is generated, animated, or sensed by computing resources and that is perceivable, editable, or manipulable by users. Such “Interactional Entities” are composable, couplable, and hybridable. “Physical” comprises matter (e.g., objects, sides, and substances), sound (e.g., audio, percussion, and noise), and any graphical item (e.g., shapes, texts, and images). For example, the sides of a cube can be considered as entities, as well as the whole cube, depending on what is sensed, mapped, or represented. Some other entities are added to this initial set that are not generated, nor animated, nor sensed by computing resources, but that support, hold, or guide entities from the initial set during interaction (e.g., surfaces, slots, and trails).
3.2. Back to Tangible User Interfaces’ Concept
According to Ullmer and Ishii: “Tangible interfaces give physical form to digital information, employing physical artifacts both as representations and controls for computational media” (Ullmer and Ishii, 2000). This definition’s phrasing reveals Ullmer and Ishii’s datum-centered viewpoint on tangible user interfaces (Shaer and Hornecker, 2010: p. 19). The adjective “tangible” should then be used to qualify data that becomes tangible, not to qualify artifacts or bodied objects (the latter is already “tangible” in the usual sense). This article claims we should then rather talk about “tangible data,” and no more about “tangible artifacts” or “tangible objects.” Thus, tangible data are data that become tangible through artifacts (that give physical form to data): only those “pure” artifacts must be qualified as tangible representations to match Ullmer and Ishii’s definition strictly. By extending this line of reasoning, this article proposes naming all artifacts according to their digital role—as instantiations of digital information—by providing meaning about how they represent digital entities. This way, the following section defines a what–how terminology to name physical representations and controls from a role-centered viewpoint.
3.3. A Role-Centered Terminology
The new terminology uses blended words (Beliaeva, 2016; Connolly, 2013) formed from a taxonomy on “what” digital entities come into the digital world and “how”. The first part concerns what digital entities are represented or controlled by physical representations:
  1. (1)
    Data refers to datum models and data states stored in computer memory and computed by the processor.
  2. (2)
    Tools are instruments that can edit the data represented in the physical world (regardless of whether data representation is tangible, graspable, or intangible).
  3. (3)
    Operations involve triggering an event or action in the computing system (e.g., core operations (Ullmer et al., 2008) include loading or saving data, loading an app, and validation).
  4. (4)
    Finally, constraints structure the behavior of data representations, thus structuring mapped digital information. They can give meaning to artifacts by triggering actions or operations (e.g., detecting a block’s presence to trigger an action (Ullmer and Ishii, 1999; Ullmer et al., 1998)). Broadly, “constraints” affect physical objects’ positioning (e.g., table surface) before being sensed by the computer. “Constraints” are not always computer-generated: for example, some circles where to place artifacts can be video-projected onto a surface or drawn with paint.
The second part concerns how digital entities come to the physical world:
  1. (1)
    Digital entities made “tangible” are represented by iconic artifacts of specialized form—with a high embodiment (of either or both representation and control).
  2. (2)
    Those that are made “graspable” are represented by symbolic artifacts of a generic form, with a low level of embodiment. Saying tangible or graspable is not saying better or worse, but simply different gradations of embodiment (Fishkin, 2004) that correspond to different design and technology answers to different needs.
  3. (3)
    As well, digital entities made “intangible” are represented by volatile computer-generated physical phenomena that cannot be grasped and disappear once the computer is turned off (e.g., display, video projection, or audio).
From the seven source words mentioned above, the seven following splinters (Barrena Jurado, 2019) are excerpted: four bases from what words: “dat-,” “tol-,” “op-,” and “const-”; and three suffixes from how words: “-ible,” “-able,” and “-nible.” Several attempts were required to select those splinters by a compromise between pronunciation, syllable count, and recognition of original words in the final terms (at least for experts aware of the terminology’s construction). Blending those bases and suffixes thus provides twelve terms embedding meaning about what and how digital entities come to the physical world (see Table 1):
  1. (1)
    Datible [”deItIb9l];
  2. (2)
    Datable [”deIt9b9l];
  3. (3)
    Datnible [”deIt.nIb9l];
  4. (4)
    Tolible [tu:lIb9l];
  5. (5)
    Tolable [tu:l9b9l];
  6. (6)
    Tolnible [tu:l.nIb9l];
  7. (7)
    Opible [OpIb9l];
  8. (8)
    Opable [Op9b9l];
  9. (9)
    Opnible [Op.nIb9l];
  10. (10)
    Constible [”kOnstIb9l];
  11. (11)
    Constable [”kOnst9b9l];
  12. (12)
    Constnible [”kOnst.nIb9l].
The first term, “datible”—“datum is tangible”—better matches the meaning of the definition of “purely tangible” user interfaces, in contrast to previous terminologies. Moreover, none of the twelve terms confuses words used for everyday objects. Finally, they provide a subtle and global descriptive power by naming four roles with three degrees in a unique namespace. The following section suggests relying on the holisticness of the terminology to characterize applicative tangible user interfaces and related physical user interfaces through vertices.
How
What Tangible Graspable Intangible
Datum DATIBLE‘Datum is tangible’ DATABLE‘Datum is graspable’ DATNIBLE‘Datum is intangible’
Tool TOLIBLE‘Tool is tangible’ TOLABLE‘Tool is graspable’ TOLNIBLE‘Tool is intangible’
Operation OPIBLE‘Operation is tangible’ OPABLE‘Operation is graspable’ OPNIBLE‘Operation is intangible’
Constraint CONSTIBLE‘Constraint is tangible’ CONSTABLE‘Constraint is graspable’ CONSTNIBLE‘Constraint is intangible’
Table 1. The twelve terms refining physical representations.
3.4. An Abstraction Through Hallmarks
[…]
To this end, we define the two following vertices:
  • Let the hallmark of an application be the vertex of entities’ count for each term, where terms are ordered first by “what” (i.e., Datum, Tangible, Operation, then Constraint), and second by “how” (i.e., Tangible, Graspable, then Intangible). The resulting hallmark is the following vertex:
    (#datible, #datable, #datnible, #tolible, #tolable, #tolnible, #opible, #opable, #opnible, #constible, #constable, #constnible).
  • Let the binary hallmark of an application be the vertex built from the hallmark where components only indicate entities’ presence for each term (i.e., components’ values greater than 1 are cut to 1).
Some distances could be calculated between these vertices. However, this article primarily proposes using hallmarks to cluster applicative tangible user interfaces into distinct classes.
3.5. Four Role-Centered Tangibility Classes
We propose to cluster applicative tangible user interfaces according to how roles are represented, with a particular focus on the role of data, whose representation is determinant to the definition of tangible user interfaces. The four resulting tangibility classes are the following:
  • In Class I specimens, data are represented only by tangible or graspable entities (i.e., presence of either datibles or datables, but absence of datnibles).
  • In Class II specimens, data are represented by intangible entities, as well as tangible or graspable entities (i.e., presence of datnibles with either datibles or datables).
  • In Class III specimens, data are represented only by intangible entities, which are coupled with the use of tangible or graspable tools (i.e., presence of datnibles combined with tolibles or tolables, but absence of datibles and datables).
  • In Class IV specimens, only operations are provided, through tangible or graspable entities (i.e., presence of opibles or opables, and data and tools are unknown or absent). This class provides specimens that are not restricted to a specific application and can thus control operations shared across applications (i.e., core operations (Ullmer et al., 2008)).
The application hallmarks’ patterns that correspond to these tangibility classes are reported in Table 2.
Hallmark pattern
What: D T O C
Class How: T G I T G I T G I T G I
I ( +, 0, 0, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$ )
( 0, +, 0, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$ )
II ( +, 0, +, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$ )
( 0, +, +, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$ )
III ( 0, 0, +, 0, +, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$ )
( 0, 0, +, +, 0, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$, $\ast$ )
IV ( 0, 0, 0, 0, 0, 0, +, 0, $\ast$, $\ast$, $\ast$, $\ast$ )
( 0, 0, 0, 0, 0, 0, 0, +, $\ast$, $\ast$, $\ast$, $\ast$ )
Note. What: ‘D’ = Data, ‘T’ = Tool, ‘O’ = Operation, ‘C’ = Constraint.
How: ‘T’ = Tangible, ‘G’ = Graspable, ‘I’ = Intangible.
‘+’ = some representations (one or more). $\ast$’ = none or some representations.
Table 2. Representations in the four tangibility classes.
The following section illustrates the descriptive power of the twelve terms through examples from the literature.
4. Illustrating on a Corpus of Applications
This section illustrates the usage of the new terminology by describing and classifying a corpus of applications from the literature.
4.1. Specimens’ Selection
Because researchers have produced hundreds of tangible user interface specimens over the past few decades (Fleck et al., 2018), the present collection is not meant to be exhaustive, but a rather representative sample. The specimen selection process is neither as rigorous as that in systematic literature reviews nor intended to provide reliable statistics. Instead, the objective is to illustrate the terminology coverage and usefulness by describing a variety of emblematic tangible user interface specimens and related physical user interfaces. Because research interests vary over the years, restricting the analysis to a specific date range could have led to the exclusion of some tangible user interface kinds; therefore, the entire research period, from the premise to the more recent work, was taken into account. The selection thus gathered various genres [^3] Seven genres of tangible user interface applications were described in 2008 (Ishii, 2008): (1) tangible telepresence, (2) tangibles with kinetic memory, (3) constructive assembly, (4) tokens and constraints, (5) interactive surfaces—tabletop TUI, (6) continuous plastic TUI, (7) augmented everyday objects, and (8) ambient media, to which actuated TUI and shape-changing TUI must now be added. of tangible user interfaces that were studied during this entire period. Finally, the selection was filtered to retain work that showed real applications. Work merely demonstrating a new technology or a new kind of tangible user interface was not retained, as its naming could vary across applications. Indeed, knowing the mapping to the digital world is necessary because being physical is not enough to be a tangible representation: the same representation can be tangible in some applications but not in others.
The selection process is summarized in Figure 1. Several sources were browsed. First, a previous state of the art (Shaer and Hornecker, 2010) and some specimens already known to the author (e.g., some specimens that were experienced during demonstration sessions at conferences over time). Second, the projects page [^4] MIT Tangible Media Group’s projects page: https://tangible.media.mit.edu/projects/, last accessed 2022-08-03. of the MIT Tangible Media Group, which is certainly the most prolific laboratory in this field (accounts for 45.5% of the final collection’s provenance). Third, to obtain even more recent work, sessions with ‘tangible’ in their title from the last ten CHI conference proceedings (from 2013 to 2022 venues). Finally, the last ten TEI conference proceedings from 2013 to 2022 were also browsed. Specimens regarding remote computer-supported conditions were rejected.
Browsing titles, abstracts,
and illustrations
Accessing sources
Extracting information from articles
Presenting a prototype?
State of the art (Shaer and Hornecker, 2010)
List of specimens that are
already known by the author
MIT TMG’s webpage
ACM CHI 2013 to 2022:
Sessions on "tangible"
ACM TEI 2013 to 2022:
Proceedings
Applicative need?
Single user or
co-presense conditions?
Genre and
subgenre
classification
Terminological
description
Hallmark
calculation
Collecting specimens
First set
Second set
Does the first set cover
the 12 terms?
Clustering by hallmarks
and sorting by genres
Third set
Yes
No
Orthogonality?
Heterogeneity?
If yes,
inclusion
Figure 1. The selection process of applicative specimens of tangible user interfaces and related physical user interfaces.:
Once many examples were added to the collection, the terms’ coverage was computed to check whether the whole terminology is relevant. Next, more examples were selected and analyzed until the full coverage of the whole terminology’s terms could be reached. Table 3 shows the number of physical representations gathered across the twelve terms.
Term Count Term Count Term Count Term Count
Datible 14 Tolible 6 Opible 8 Constible 3
Datable 11 Tolable 17 Opable 12 Constable 19
Datnible 38 Tolnible 8 Opnible 7 Constnible 2
Table 3. Collection’s physical representation count by terms.
The orthogonality of the specimens across the terminology was also checked to assess the collection’s diversity. To this end, the hallmarks and binary hallmarks of the applications were computed.
4.2. Corpus Description
Tables 4, 5, and 6 show the resulting naming of the 145 physical entities from the collection of applications.
The four roles enabled sorting all of these physical entities. Data is the most prevalent role, accounting for 43% of the physical entities. However, 57% of the physical entities are not data: 21% are tools, 19% are operations, and 17% are constraints. Therefore, overemphasizing design or research on data representations would overlook a significant portion of the entities encountered in applicative tangible user interfaces. Furthermore, the terminology enabled the naming of all of the physical entities found in the 33 applications. Being able to distinguish and designate entities might be useful to designers. For example, six of the applications are depicted in Figure 2 with their what–how naming.
Application Physical Representation
What How
# Name Year References Name or Description D T O C T G I WHT Name
1. Slot Machine 1976 (Perlman, 1976) and see Rows (procedure) - - - $\times$ - $\times$ - CONSTABLE
in (Ullmer et al., 2022: Ch. 1) Buttons (on each row) - - $\times$ - - $\times$ - OPABLE
Plastic cards (commands) - - $\times$ - - $\times$ - OPABLE
Display triangle (turtle) $\times$ - - - - - $\times$ DATNIBLE
Circular robot (turtle) $\times$ - - - - $\times$ - DATABLE
2. CAAD 3D Modelling System 1979 (Aish, 1979; Aish and Noakes, 1984) Building Blocks $\times$ - - - - $\times$ - DATABLE
Design geometry (perspective view) $\times$ - - - - - $\times$ DATNIBLE
Evaluation measures (Isoplots) $\times$ - - - - - $\times$ DATNIBLE
3. Self-Builder Model (Segal Model) 1980 (Frazer and Frazer, 1982, 1980) cited Panels $\times$ - - - $\times$ - - DATIBLE
from (Sutphen et al., 2000) Board - - - $\times$ - $\times$ - CONSTABLE
Wireframe rendering $\times$ - - - - - $\times$ DATNIBLE
Feedback tool (house area, cost) $\times$ - - - - - $\times$ DATNIBLE
4. Marble Answering Machine 1992 See in Follmer et al., 2013, Ishii and Ullmer,1997 Machine - - - $\times$ - $\times$ - CONSTABLE
and (Ullmer et al., 2022: Ch. 1) Machine’s indentation (play slot) - - $\times$ - - $\times$ - OPABLE
Marble (message) $\times$ - - - $\times$ - - DATIBLE
Message $\times$ - - - - - $\times$ DATNIBLE
5. Head Prop 1994 (Hinckley et al., 1994) Doll’s head (brain) $\times$ - - - $\times$ - - DATIBLE
Plate (slicing) - $\times$ - - $\times$ - - TOLIBLE
Plate thumb button (clutch) - - $\times$ - - $\times$ - OPABLE
Foot pedal (clutch) - - $\times$ - - $\times$ - OPABLE
Brain 2D slice $\times$ - - - - - $\times$ DATNIBLE
6. GraspDraw 1995 (Fitzmaurice et al., 1995) Bricks - $\times$ - - - $\times$ - TOLABLE
2D shapes $\times$ - - - - - $\times$ DATNIBLE
Inkwell - - $\times$ - $\times$ - - OPIBLE
Functions’ tray (select, delete, shapes) - - $\times$ - - $\times$ - OPABLE
ActiveDesk surface - - - $\times$ - $\times$ - CONSTABLE
7. MetaDESK 1997 (Ishii and Ullmer, 1997) Phicons $\times$ - - - $\times$ - - DATIBLE
Campus Map $\times$ - - - - - $\times$ DATNIBLE
Overlay view (passive lens) - - $\times$ - $\times$ - - OPIBLE
3D view (active lens) $\times$ - - - - - $\times$ DATNIBLE
Scaling and Rotating Device - $\times$ - - - $\times$ - TOLABLE
Desk’s surface - - - $\times$ - $\times$ - CONSTABLE
8. Build-IT 1997 (Rauterberg et al., 1997, 1998; Fjeld et al., 1999) Bricks - $\times$ - - - $\times$ - TOLABLE
Plan view $\times$ - - - - - $\times$ DATNIBLE
Objects (robots, tables…) $\times$ - - - - - $\times$ DATNIBLE
Virtual cameras - - $\times$ - - - $\times$ OPNIBLE
EyeCatchers - - $\times$ - - - $\times$ OPNIBLE
Table surface - - - $\times$ - $\times$ - CONSTABLE
3D view $\times$ - - - - - $\times$ DATNIBLE
9. Pinwheels 1998 (Wisneski et al., 1998) Pinwheels $\times$ - - - $\times$ - - DATIBLE
10. Voodoo Dolls 1998 (Isidoro and Sclaroff, 1998) Spongy object $\times$ - - - - $\times$ - DATABLE
Graphics model $\times$ - - - - - $\times$ DATNIBLE
11. mediaBlocks 1998 (Ullmer and Ishii, 1999; Ullmer et al., 1998) mediaBlock - - $\times$ - - $\times$ - OPABLE
Slots - - - $\times$ - $\times$ - CONSTABLE
12. musicBottles 1999 (Ishii et al., 1999, 2001a) Music $\times$ - - - - - $\times$ DATNIBLE
Bottle (music) $\times$ - - - - $\times$ - DATABLE
Cork - - $\times$ - $\times$ - - OPIBLE
Triangular table - - - $\times$ - $\times$ - CONSTABLE
Central "stage" area - - - $\times$ - - $\times$ CONSTNIBLE
13. Urp (Urban Planning Workbench) 1999 (Ben-Joseph et al., 2001; Ishii et al., 2002; Underkoffler and Ishii, 1999) Architectural Model $\times$ - - - $\times$ - - DATIBLE
Road-object (strips) $\times$ - - - $\times$ - - DATIBLE
Material-transformation-object (wand) - $\times$ - - - $\times$ - TOLABLE
Video-camera-object - $\times$ - - $\times$ - - TOLIBLE
Clock-object - $\times$ - - $\times$ - - TOLIBLE
Wind-generating tool - $\times$ - - - $\times$ - TOLABLE
Anemometer-object (arrow) - - $\times$ - - $\times$ - OPABLE
Wind magnitude (number) - - $\times$ - - - $\times$ OPNIBLE
Distance-measuring-object - - $\times$ - - $\times$ - OPABLE
Distance (number) - - $\times$ - - - $\times$ OPNIBLE
Shadows/Relections $\times$ - - - - - $\times$ DATNIBLE
Airflow grid $\times$ - - - - - $\times$ DATNIBLE
Workbench - - - $\times$ - $\times$ - CONSTABLE
14. Senseboard 2001 (Jacob et al., 2002) Grid - - - $\times$ - - $\times$ CONSTNIBLE
Rectangular pucks $\times$ - - - - $\times$ - DATABLE
View detail puck - $\times$ - - - $\times$ - TOLABLE
Arrow puck - $\times$ - - - $\times$ - TOLABLE
Values $\times$ - - - - - $\times$ DATNIBLE
Board - - - $\times$ - $\times$ - CONSTABLE
Note. “-” = empty. What: ‘D’ = Data, ‘T’ = Tool, ‘O’ = Operation, ‘C’ = Constraint. How: ‘T’ = Tangible, ‘G’ = Graspable, ‘I’ = Intangible. ‘WHT Name’ = What-How Terminology name.
Table 4. Physical representations named for 14 applications from the literature, ordered by ascending years, from 1976 to 2001.
Application Physical Representation
What How
# Name Year References Name or Description D T O C T G I WHT Name
15. Illuminating Clay 2002 (Ishii et al., 2004; Piper et al., 2002) Clay Model $\times$ - - - $\times$ - - DATIBLE
Rotative plaform - - - $\times$ - $\times$ - CONSTABLE
Crosshairs - $\times$ - - - - $\times$ TOLNIBLE
Cross Sections $\times$ - - - - - $\times$ DATNIBLE
Analysis Function Thumbnails $\times$ - - - - - $\times$ DATNIBLE
3-D perspective view $\times$ - - - - - $\times$ DATNIBLE
16. AudioPad 2002 (Patten et al., 2002, 2006) Sounds $\times$ - - - - - $\times$ DATNIBLE
Pucks (audio tracks) $\times$ - - - - $\times$ - DATABLE
Star-shape puck (sound selector) - $\times$ - - - $\times$ - TOLABLE
Hierarchical menu - - $\times$ - - - $\times$ OPNIBLE
SenseTable surface - - - $\times$ - $\times$ - CONSTABLE
17. ReacTable 2003 (Jordà et al., 2007) Music $\times$ - - - - - $\times$ DATNIBLE
Point (audio output) $\times$ - - - - - $\times$ DATNIBLE
Square puck (audio source) $\times$ - - - - $\times$ - DATABLE
Rounded square puck (filter) - $\times$ - - - $\times$ - TOLABLE
Round puck (controller) - $\times$ - - - $\times$ - TOLABLE
Decagon puck (control filter) - $\times$ - - - $\times$ - TOLABLE
Pentagon puck (audio mixer) - $\times$ - - - $\times$ - TOLABLE
Lines (audio flow) $\times$ - - - - - $\times$ DATNIBLE
Table surface - - - $\times$ - $\times$ - CONSTABLE
18. IP Network Design Workbench 2003 (Kobayashi et al., 2003) Pucks - $\times$ - - - $\times$ - TOLABLE
Nodes $\times$ - - - - - $\times$ DATNIBLE
Links $\times$ - - - - - $\times$ DATNIBLE
Nodes menu - - $\times$ - - - $\times$ OPNIBLE
Parameter puck (with button) - $\times$ - - - $\times$ - TOLABLE
Link bandwidth - $\times$ - - - - $\times$ TOLNIBLE
Router service priority - $\times$ - - - - $\times$ TOLNIBLE
Number of client users - $\times$ - - - - $\times$ TOLNIBLE
Server performance - $\times$ - - - - $\times$ TOLNIBLE
Table surface - - - $\times$ - $\times$ - CONSTABLE
Measurement graphs - - $\times$ - - - $\times$ OPNIBLE
19. Query Shapes 2004 (Ichida et al., 2004) Cubes (ActiveCubes) $\times$ - - - - $\times$ - DATABLE
Voxel data representation $\times$ - - - - - $\times$ DATNIBLE
3D shape models $\times$ - - - - - $\times$ DATNIBLE
20. TUISTER 2004 (Butz et al., 2004) TUISTER - - $\times$ - $\times$ - - OPIBLE
21. I/O Brush 2004 (Ryokai et al., 2004, 2007) Brush - $\times$ - - $\times$ - - TOLIBLE
World (palette) - - $\times$ - $\times$ - - OPIBLE
Canvas $\times$ - - - - - $\times$ DATNIBLE
22. PICO 2005 (Patten and Ishii, 2007) Pucks (antenas) $\times$ - - - - $\times$ - DATABLE
Map (city) $\times$ - - - - - $\times$ DATNIBLE
Flexible curve - - - $\times$ $\times$ - - CONSTIBLE
Rubber band - - - $\times$ $\times$ - - CONSTIBLE
Collar - - - $\times$ $\times$ - - CONSTIBLE
Teflon / Sandpaper - - - $\times$ - $\times$ - CONSTABLE
Table surface - - - $\times$ - $\times$ - CONSTABLE
23. ArcheoTUI 2007 (Reuter et al., 2007) Props $\times$ - - - - $\times$ - DATABLE
3D fragments $\times$ - - - - - $\times$ DATNIBLE
Foot pedals (clutch) - - $\times$ - - $\times$ - OPABLE
24. Slurp 2007 (Zigelbaum et al., 2008) Eyedropper - - $\times$ - $\times$ - - OPIBLE
25. GeoTUI 2008 (Couture et al., 2008) Map (cube top view) $\times$ - - - - - $\times$ DATNIBLE
Cutting line - $\times$ - - - - $\times$ TOLNIBLE
Two-puck prop - $\times$ - - - $\times$ - TOLABLE
Ruler prop - $\times$ - - $\times$ - - TOLIBLE
Button box - - $\times$ - - $\times$ - OPABLE
Tabletop - - - $\times$ - $\times$ - CONSTABLE
26. Relief 2010 (Leithinger and Ishii, 2010; Leithinger et al., 2011) 2.5D shape display (terrain) $\times$ - - - $\times$ - - DATIBLE
Topographical map $\times$ - - - - - $\times$ DATNIBLE
27. Teegi 2014 (Frey et al., 2014) Teegi character (user’s activity) $\times$ - - - $\times$ - - DATIBLE
Brain model (user’s activity) $\times$ - - - $\times$ - - DATIBLE
Filter area - $\times$ - - - - $\times$ TOLNIBLE
Mini-Teegis (filters) - $\times$ - - $\times$ - - TOLIBLE
EEG Raw Signals $\times$ - - - - - $\times$ DATNIBLE
Color map amplitude - $\times$ - - - - $\times$ TOLNIBLE
Color map cursor - $\times$ - - - $\times$ - TOLABLE
Table surface - - - $\times$ - $\times$ - CONSTABLE
28. SoundFORMS 2016 (Colter et al., 2016) Trigger pins - - $\times$ - - $\times$ - OPABLE
Soundwave pins $\times$ - - - $\times$ - - DATIBLE
Sound $\times$ - - - - - $\times$ DATNIBLE
Note. “-” = empty. What: ‘D’ = Data, ‘T’ = Tool, ‘O’ = Operation, ‘C’ = Constraint. How: ‘T’ = Tangible, ‘G’ = Graspable, ‘I’ = Intangible. ‘WHT Name’ = What-How Terminology name.
Table 5. Physical representations named for 14 applications from the literature, ordered by ascending years, from 2002 to 2016.
Application Physical Representation
What How
# Name Year References Name or Description D T O C T G I WHT Name
29. reSpire 2019 (Choi et al., 2019) Shape-changing fabric $\times$ - - - $\times$ - - DATIBLE
30. CairnFORM 2019 (Daniel et al., 2019) Ring chart $\times$ - - - - $\times$ - DATABLE
31. Embodied Axes 2020 (Cordeil et al., 2020) Orthogonal arms (data axes) - - $\times$ - $\times$ - - OPIBLE
32. CoDa 2020 (Veldhuis et al., 2020) Tokens (data points) $\times$ - - - $\times$ - - DATIBLE
Interactive surface - - - $\times$ - $\times$ - CONSTABLE
Sidebar buttons (filter and analytic functions) - $\times$ - - - $\times$ - TOLABLE
Data points $\times$ - - - - - $\times$ DATNIBLE
Analytical functions (lines) $\times$ - - - - - $\times$ DATNIBLE
33. SABLIER 2022 (Mahieux et al., 2022) Hourglass - - $\times$ - $\times$ - - OPIBLE
Note. “-” = empty. What: ‘D’ = Data, ‘T’ = Tool, ‘O’ = Operation, ‘C’ = Constraint. How: ‘T’ = Tangible, ‘G’ = Graspable, ‘I’ = Intangible. ‘WHT Name’ = What-How Terminology name.
Table 6. Physical representations named for 5 applications from the literature, ordered by ascending years, from 2019 to 2022.
(a) Relief: Terrain dynamic exploration (#26).:
(b) SABLIER: Time navigation in cultural heritage scopes (#33).:
(c) Head Prop: Neurosurgical visualization (#5).:
(d) ArcheoTUI: Virtual archaeology fragments’ assembly (#23).:
(e) GraspDraw: 2D shape edition (#6).:
(f) GeoTUI: Cutting line selection for subsoil exploration (#25).:
Licenses: (b) Castle 3D model: Ramon Duran, CC0; (c) Doll’s head 3D model: David Gilman, $C$ BY 4.0 – Brain MRI picture: Rudy Liggett, CC0; (d) Vase 3D model: Robert Tolliver, $C$ BY 4.0;(f) Topographic map: Gordon Dylan Johnson, CC0; (c,d,e,f) Chair 3D model: Zach Steindler, MIT License.
Figure 2. Pointing out physical representations and controls for six depicted example specimen applications.
The hallmarks of the 33 applicative tangible user interfaces are reported in Table 7. This collection provides 28 orthogonal hallmarks. Indeed, 8 applications that share the same hallmarks are distributed into 3 clusters [^5] The three clusters’ hallmarks are and (1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0): Relief (#26), and reSpire (#29); (0, 1, 2, 0, 0, 0, 0, 0, 0, 0, 0, 0): CAAD 3D Modelling System (#2) and Query Shapes (#19); (0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0): TUISTER (#20), Slurp (#24), Embodied Axes (#31), and SABLIER (#33). . The specimens in these clusters are of various tangible user interface genres. However, some specimens of the same genre (e.g., AudioPad (#16) and ReacTable (#17)) have different hallmarks, thus demonstrating that the new holistic terminology is able to discriminate physical user interfaces at the application level (i.e., beyond their morphological and technological descriptions). This collection also provides 26 orthogonal binary hallmarks (i.e., 10 applications are now distributed into the 3 clusters [^6] Pinwheels (#9) joins the cluster of Relief (#26) and reSpire (#29) because sharing the same binary hallmark: (1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0). Voodoo Dolls (#10) joins the cluster of CAAD 3D Modelling System (#2) and Query Shapes (#19) because having the same binary hallmark: (0, 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0). ).
# Application Hallmark Class
1. Slot Machine (0, 1, 1, 0, 0, 0, 0, 2, 0, 0, 1, 0) II
2. CAAD 3D Modelling System (0, 1, 2, 0, 0, 0, 0, 0, 0, 0, 0, 0) II
3. Self-Builder Model (Segal Model) (1, 0, 2, 0, 0, 0, 0, 0, 0, 0, 1, 0) II
4. Marble Answering Machine (1, 0, 1, 0, 0, 0, 0, 1, 0, 0, 1, 0) II
5. Head Prop (1, 0, 1, 1, 0, 0, 0, 2, 0, 0, 0, 0) II
6. GraspDraw (0, 0, 1, 0, 1, 0, 1, 1, 0, 0, 1, 0) III
7. MetaDESK (1, 0, 2, 0, 1, 0, 1, 0, 0, 0, 1, 0) II
8. Build-IT (0, 0, 3, 0, 1, 0, 0, 0, 2, 0, 1, 0) III
9. Pinwheels (N, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0) I
10. Voodoo Dolls (0, 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0) II
11. mediaBlocks (0, 0, 0, 0, 0, 0, 0, 1, 0, 0, 1, 0) IV
12. musicBottles (0, 1, 1, 0, 0, 0, 1, 0, 0, 0, 1, 1) II
13. Urp (Urban Planning Workbench) (2, 0, 2, 2, 2, 0, 0, 2, 2, 0, 1, 0) II
14. Senseboard (0, 1, 1, 0, 2, 0, 0, 0, 0, 0, 1, 1) II
15. Illuminating Clay (1, 0, 3, 0, 0, 1, 0, 0, 0, 0, 1, 0) II
16. AudioPad (0, 1, 1, 0, 1, 0, 0, 0, 1, 0, 1, 0) II
17. ReacTable (0, 1, 3, 0, 4, 0, 0, 0, 0, 0, 1, 0) II
18. IP Network Design Workbench (0, 0, 2, 0, 2, 4, 0, 0, 2, 0, 1, 0) III
19. Query Shapes (0, 1, 2, 0, 0, 0, 0, 0, 0, 0, 0, 0) II
20. TUISTER (0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0) IV
21. I/O Brush (0, 0, 1, 1, 0, 0, 1, 0, 0, 0, 0, 0) III
22. PICO (0, 1, 1, 0, 0, 0, 0, 0, 0, 3, 2, 0) II
23. ArcheoTUI (0, 1, 1, 0, 0, 0, 0, 1, 0, 0, 0, 0) II
24. Slurp (0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0) IV
25. GeoTUI (0, 0, 1, 1, 1, 1, 0, 1, 0, 0, 1, 0) III
26. Relief (1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0) I
27. Teegi (2, 0, 1, 1, 1, 2, 0, 0, 0, 0, 1, 0) II
28. SoundFORMS (1, 0, 1, 0, 0, 0, 0, 1, 0, 0, 0, 0) II
29. reSpire (1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0) I
30. CairnFORM (0, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0) I
31. Embodied Axes (0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0) IV
32. CoDa (1, 0, 2, 0, 1, 0, 0, 0, 0, 0, 1, 0) II
33. SABLIER (0, 0, 0, 0, 0, 0, 1, 0, 0, 0, 0, 0) IV
Table 7. Hallmarks and classes of the 33 applications.
4.3. Corpus Classification
The classification of the corpus through the four classes is shown in Table 7: the collection comprises 4 applications of Class I, 19 of Class II, 5 of Class III, and 5 of Class IV. All of the applications could be distributed among only the four classes. Only a few applications of the corpus represented data through only tangible or graspable representation (i.e., 4 applications of Class I), whereas over half combined them with intangible data representations (i.e., 19 applications of Class II). Moreover, only a few applications merely combine intangible data representations with tangible or graspable tools (5 applications of Class III). However, this distribution may result from the corpus’s constitution (i.e., by selecting orthogonal hallmarks), and may differ in a more extended corpus (e.g., by gathering applications through a systematic literature review).
The collection shows several possible configurations for Class I applications, from duplication of the same datibles, to datables combined with tolables and opnibles, to a unique datible. For example, the ambient interface Pinwheels (#9) uses multiple replicated datibles: the “pinwheels” rotation speed then informs about a flow (e.g., stock market exchange, car traffic, or e-mail traffic). CairnFORM (#30) is one datible made of replicated rings, whose expansion informs about renewable energy forecasts. Finally, Relief (#26) and reSpire (#29) are single deformable datibles, that change surfaces’ shapes to modelize terrain topography and echo self-breathing, respectively.
The collection includes cases for Class II applications where datnibles are bound or combined with datibles or datables. For example, the Head Prop specimen (#5) binds the intersection between a datible (“doll’s head”) and a tolible (“plate”) to a datnible (“slice view”). On the tabletop ReacTable (#17), audio sources are represented by datables (“square pucks”), audio flow and audio output are materialized by datnibles (some “lines” and a “dot,” respectively); some tolables allow mixing or filtering the audio flow. Finally, the actuated interface SoundFORMS (#28) assigns digital roles to an array of actuated pins: pressing some opables (“trigger pins”) starts playing datnibles (“sounds”) that are bound to datibles (“soundwave pins”).
In Class III applications, the tangible or graspable part of the user interface is not about data, but rather tools or operations. For example, the I/O Brush (#21) uses World as an opible (“palette”), to draw with a tolible (“brush”) on a datnible (“canvas”). The tabletop GraspDraw (#6) edits datnibles (“shapes”) using a tolable (“pucks”) that serve as physical anchors. Finally, the tabletop GeoTUI (#25) controls a tolnible (“cutting line”) with a tolible (“ruler”) or a tolable (“two-puck prop”) on a datnible (“top view map”).
Artifacts of Class IV are designed to integrate across various applications, whose scopes can extend beyond tangible user interfaces (e.g., graphical user interfaces (Zigelbaum et al., 2008), augmented reality (Cordeil et al., 2020), and virtual reality caves (Mahieux et al., 2022)). For example, Slurp (#24) is an “eyedropper” that can extract and inject digital information. TUISTER (#20), Embodied Axes (#31), and SABLIER (#33) provide artifacts to browse hierarchical structures, control 3D positions, and navigate through space and time, respectively.
5. Mapping the Previous Terms
Figure 3 illustrates a mapping between the previous terms found in the literature and the new holistic terminology, achieved through a projection onto the two axes of the taxonomy that originated the terminology (i.e., “what” is physically represented and “how”). Only one term can cover the whole terminology. Except for this holistic term, the mapping reveals a focus of previous research on representing tangible data, which is describable through half of the terms. In contrast, constraints and intangible data are describable by only one term each, and intangible tools and operations are outside the conceptual perception.
ASUR++ Terms (Dubois et al., 2003)
General Terms
Examples of Media
Legend:
TAC Terms (Shaer et al., 2004)
Isolated Terms
TANGIBLE
GRASPABLE
INTANGIBLE
TOOL
OPERATION
CONSTRAINT
HOW
WHAT
Core Tangibles (Ullmer et al., 2008)
Pyfos (Shaer et al., 2004)
Constraints
(Shaer et al., 2004; Ullmer, 2002; Ullmer et al., 2005)
DATUM
Intangible representations (Ullmer, 2002; Ullmer et al., 2005)
Tangible representations (Ullmer, 2002; Ullmer et al., 2005)
Robject (Dubois et al., 2003)
Sound
Rtool (Dubois et al., 2003)
Video
Passive Props (Hinckley et al., 1994)
Phicons (Ullmer and Ishii, 1997)
Artifacts (Ullmer and Ishii, 2000)
Tangibles (Ullmer and Ishii, 2000)
Tokens
(Shaer et al., 2004; Ullmer, 2002; Ullmer et al., 2005)
Phandles (Ullmer and Ishii, 1997)
Phidgets (Greenberg and Fitchett, 2001)
Figure 3. Mapping thirteen previous terms for physical representations and controls on the new conceptual space.:
The distorted focus on naming tangible data may have its origins in the community’s search for paradigms and frameworks, as well as the influence of the data-centered viewpoint in the definition of tangible user interfaces. However, other roles than data are required to describe applicative tangible user interfaces.
The following section proposes to match the four tangibility classes with some previous classifications through the set of (orthogonal) applications provided by the corpus.
6. Cross-Matching Classifications
Some previous classifications have discriminated between tangible user interfaces according to their morphology (i.e., how they are physically shaped, such as surfaces (Ishii, 2008), everyday objects (Ishii, 2008), or discrete tangibles (Ishii et al., 2012)) and behavior (i.e., how users interact with them and how they interact with users, such as assembly (Ishii, 2008), deformation (Ishii et al., 2012), or transformation (Ishii et al., 2012)). Figure 4 bridges these previous classifications and the tangibility classes through the collection of applicative tangible user interfaces.
Matching previous classifications, the applications’ tangibility classes are split into five clusters. The present collection of applications already shows that previous classifications and tangibility classes have no direct correspondence. Whereas “Artifacts & Objects” gathers applications coming from the four classes, “Ambient Media” and “Constructive Assemblies” count only one class each. Moreover, this initial matching may evolve when more applications are included. Indeed, even if some genres fit implicitly into certain classes well, based on their definition (e.g., interactive surfaces, such as tabletops, falign well with Class II and Class III), some classes may appear in certain clusters. For example, constructive assemblies could also be stackable audio or video filters that are tools controlling intangible data (i.e., describing such an application specimen would then require bridging Class III with “Constructive Assemblies”). Furthermore, other clusters may appear because certain genres of tangible user interfaces are absent from the collection of applications (e.g., “Tangible Telepresence” (Ishii, 2008) and “Tangibles with Kinetic Memory” (Ishii, 2008)). However, we believe that the descriptive power of the four classes may provide a sufficient design space to integrate any new applicative tangible user interface and related physical user interfaces.
Subgenre
Genre
Tangibility Class
Application
Dynamic everyday objects
(#9) Pinwheels
Ambient Media
Class I
Tokens and
Constraints
Interactive Surfaces
and Spaces
(#22) PICO
(#14) Senseboard
(#7) MetaDESK
(#16) AudioPad
(#17) ReacTable
(#32) CoDa
(#6) GraspDraw
(#8) Build-IT
(#18) IP Network Design Workbench
(#25) GeoTUI
Actuated tabletop
Board
Tabletop
Workbench
(#13) Urp
(#1) Slot Machine
(#4) Marble Answering Machine
(#11) mediaBlocks
Class II
Class III
Class II
Class IV
Artifacts & Objects
Constructive
Assemblies
Mirrored constructive assemblies
(#2) CAAD 3D Modelling System
(#3) Self-Builder Model
(#19) Query Shapes
(#21) I/O Brush
(#20) TUISTER
(#31) Embodied Axes
(#33) SABLIER
Class II
Class III
Class II
Class IV
Symbolic tokens (Ullmer and Ishii, 2000)
Class I
Transformable continuous tangibles (Ishii et al., 2012)
(#29) reSpire
(#30) CairnFORM
(#15) Illuminating Clay
Deformable continuous tangibles (Ishii et al., 2012)
Augmented everyday objects
(#27) Teegi
(#24) Slurp
(#12) musicBottles
(#26) Relief
Mirrored manipulative artifacts
(#5) Head Prop
(#10) Voodoo Dolls
(#23) ArcheoTUI
(#28) SoundFORMS
Mirrored everyday objects
Manipulative artifacts
(Ishii, 2008)
(Ishii, 2008)
(Ishii, 2008)
(Ishii, 2008)
(Ishii, 2008)
(Ishii, 2008)
Figure 4. Matching previous classifications with what–how tangibility classes through a collection of applications.:
7. Limitations and Future Work
A limitation of this work is that analyses were not consolidated with the specimens’ designers. Describing the 33 applications was mostly straightforward, but sometimes required advanced questioning. An example is when describing the Pinwheels specimen (#9), which represents information flow in an ambient manner. At first glance, Pinwheels must be datnible: airflow represents the datum flow. However, this idea evolved from using airflow in the ambientROOM (Ishii et al., 1998)—merely needing air coming into the room through a hole—to using visual “spinning pinwheels” (Dahley et al., 1998; Ishii et al., 2001b; Wisneski et al., 1998). Hence, Pinwheels are described as datibles [^7] Treating single datnibles (airflow) as physical user interfaces would lead to reconsidering Class I definition or to creating a new class.. Another example lies in the vertical measurement graphs (i.e., usage rate, response time, and running cost graphs) placed next to the IP Network Design Workbench (#18). Some could have considered these graphs as auxiliary GUIs. However, they were considered as opnibles (i.e., intangible measurement operations), as well as measurements in the Urp (#13) that were considered as operations. Advanced questioning can become tricky when design thoughts are not clearly detailed in the articles. The community must benefit from future work that describes previous physical user interface specimens in a publicly shared database (e.g., (Fleck et al., 2018)) and future articles that describe specimens using holistic terminologies.
This work focused on the roles found in tangible and related physical user interfaces without considering remote collaboration conditions (Brave et al., 1998). However, in such conditions, remote users may be represented by the user interface through streams (Ishii, 1990; Ishii and Kobayashi, 1992; Leithinger et al., 2014), shapes (Brave and Dahley, 1997; Brewer et al., 2007), or avatars (Lee et al., 2025) and even interact with physical entities (Brave and Dahley, 1997; Leithinger et al., 2014). Therefore, these conditions raise the question of defining a role for remote users specifically (or considering them merely as data or tools?), refining their body parts (e.g., head, forearms, and hands), and defining terms that describe how they appear in the user interface. These questions may extend the terminology to interactions with robots and cobots in Human-Robot Interaction (Stoeva et al., 2024) and Extended Reality conditions (Gonzalez-Sanchez and Guerra-Silva, 2025).
The description of bodied entities is refined on only two levels of embodiment (i.e., tangible and graspable), which roughly simplifies the twenty values of Fishkin’ taxonomy (Fishkin, 2004). However, this article considered a level of embodiment for bodiless entities (i.e., intangible), even if intangible entities can also encounter various levels of representativeness (e.g., a beep or slang audio feedback when sending a file to trash, versus paper-creasing noise).
This work separated hallmarks into classes. However, defining a distance calculation between hallmarks could serve the needs of automatic classification algorithms (e.g., cladistics (Fleck et al., 2018)) that would help sort, organize, and browse the vast collection of specimens provided by the field of tangible user interfaces over the past three decades.
This article introduces the concept of interactional entities briefly, which future work will need to define more extensively.
8. Conclusion
This article introduces a “what–how” terminology (WHT) that provides twelve terms to name physical representations and controls: by taking a role-centered viewpoint (i.e., the digital role), the resulting terms self-contain meaning about how digital entities come into the physical world. Those terms are blended words, whose logical structure from known words must make them easy to learn, recollect, and understand. Mainly, this new terminology outperforms the previous terms found in the literature for three reasons. First, it defines all terms within a single namespace, free from specific paradigms, thus allowing for cross-genre descriptions. Second, it refines the descriptive level of physical user interface components. Third, it better aligns with definitions, thereby enhancing communication and understanding. This terminology introduces the term “datible,” which matches the definition of pure tangible user interfaces, along with providing eleven terms that distinguish other representations. This terminology must benefit at least three scopes: first, supporting ideation by providing vocabulary that describes possibilities; second, offering a holistic conceptual space for toolkit architectures; and finally, offering a taxonomy usable for the exploration and classification of physical user interfaces.
Acknowledgements.
The author thanks the reviewers of the ACM CHI 2023 conference for their valuable comments, which helped revise and improve certain sections of this article, as well as extend its contributions.